# Inverse Reinforcement Learning with Switching Rewards and History Dependency for Characterizing Animal Behaviors

### v1

@@ 1 Introduction
Inverse reinforcement learning (IRL), which infers agents’ policies and intrinsic reward functions based on their interactions with the environment (Ng & Russell, 2000; Abbeel & Ng, 2004; Ziebart et al., 2008; 2010; Wu et al., 2024), has been shown to be effective in capturing animal decision-making intentions by learning reward functions from behavioral trajectories (Sezener et al., 2014; Yamaguchi et al., 2018; Pinsler et al., 2018; Hirakawa et al., 2018). However, traditional IRL assumes a single static reward function over time, limiting its ability to account for shifts in intrinsic motivations. To address this limitation, recent IRL variants have aimed to uncover heterogeneous and time-varying reward functions (Babes-Vroman et al., 2011; Surana & Srivastava, 2014; Nguyen et al., 2015; Ashwood et al., 2022a; Zhu et al., 2024). Despite these advancements, a significant challenge remains unaddressed: animals make decisions based on their history, not just their current state (Kennedy, 2022; Hattori et al., 2019). For example, in perceptual decision-making tasks, mice are found to make new decision based on reward, state and decision history (Ashwood et al., 2022b). Incorporating historical context into the modeling could offer a more accurate representation of animal behavior.
@@
To address the absence of history dependency in time-varying IRL models, we introduce a novel framework called SWitching IRL (SWIRL). Similar to Zhu et al. (2024), SWIRL models long recordings of animal behaviors as a sequence of short-term decision-making processes. Each decision-making process is treated as a Markov decision process (MDP) with a unique reward function that can be inferred using IRL. The segmentation of a long recording into switching decision-making processes is unknown; therefore, each process is regarded as being associated with a hidden mode that must also be inferred. Most importantly, SWIRL incorporates biologically plausible history dependency, drawing on insights from animal behavior. The history dependency is added at two levels: the transitions between decision-making processes (decision-level) and the actions taken to achieve a single goal within a decision-making process (action-level). Decision-level dependency is reflected in the transitions between decision-making processes over extended sequences of time bins, suggesting that an animal’s current choice is shaped by its previous decisions and environmental feedback. Additionally, we posit that these transitions are influenced by the animal’s location. For example, after a mouse drinks from a water port, if it stays nearby, it is more likely to seek another goal. Conversely, if it is far from the port, indicating it has been away for some time, it may become thirsty again and return to search for water. For action-level history dependency, we will model the policy and reward functions as dependent on trajectory history within each decision-making process, using a non-Markovian decision framework. Such a dependency has been studied by existing reinforcement learning research, which often characterizes exploration with reward functions based on historical states and actions (Houthooft et al., 2016; Sharafeldin et al., 2024). Importantly, our paper is the first to incorporate history-dependent policies and rewards into IRL.
@@ 2 Related Work
IRL for animal behavior understanding. IRL has been widely used to infer animals’ behavioral strategies and decision-making policies when the reward is unknown. For instance, Pinsler et al. (2018) applies IRL to uncover the unknown reward functions of pigeons, explaining and reproducing their flock behavior, and developed a method to learn a leader-follower hierarchy. Similarly, Hirakawa et al. (2018) uses IRL to learn reward functions from animal trajectories, identifying environmental features preferred by shearwaters, and discovered differences in male and female migration route preferences based on the estimated rewards. In another study, Yamaguchi et al. (2018) apllies IRL to C. elegans thermotactic behavior, revealing distinct behavioral strategies for fed and unfed worms. Additionally, Sezener et al. (2014) maps reward functions for rats freely moving in a square area, showing how these rewards changed before and after training. While these studies demonstrate the utility of IRL in uncovering behavioral strategies of freely moving animals, they share a key limitation: they all assume a single reward function governs all animal behaviors, which does not account for the complexities of long-term decision-making.
@@
Dynamics-based behavior analysis in animal neuroscience. Traditional approaches to analyzing animal behavior in neuroscience often rely on autoregressive dynamics models. For instance, MoSeq and related works (Wiltschko et al., 2015; Weinreb et al., 2024) assume that animal behavior consists of multiple segments modeled by an HMM, with each segment evolving through an autoregressive process. Stone (2023) introduces a switching linear dynamical system (SLDS), similar to an AR-HMM, but with an additional layer of continuous latent states between the behavioral trajectories and the hidden states representing behavioral segments. We argue that if each segment lasts only a few seconds, it represents meaningful action motifs, such as grooming and sniffing. However, if a segment is significantly longer and reflects a decision-making process, traditional dynamics-based models may not be suitable for identifying these long-term segments. However, these dynamics-based models are not entirely independent of SWIRL. We will demonstrate that SWIRL generalizes purely dynamics-based models by relying on a more principled IRL framework to identify multiple decision-making processes. Our goal is to offer profound insights that bridge these traditional and new models for animal behavioral analysis.
@@ 3.2 Inverse Reinforcement Learning
Inverse Reinforcement Learning (IRL) addresses the scenario where we have gathered multiple trajectories from an expert agent $\pi^{*}$, comprising a set of state-action pairs $\{(s^{*}_{t},a^{*}_{t})\}$. The goal is to estimate the policy and reward that generated these state-action pairs, often referred to as demonstrations in the literature. We assume that we have collected $N$ expert trajectories, denoted as $\mathcal{D}=\{\xi_{1},\xi_{2},\ldots,\xi_{N}\}$. Each trajectory consists of a sequence of state-action pairs, represented as $\xi_{n}=\{(s_{1}^{*},a_{1}^{*}),(s_{2}^{*},a_{2}^{*}),\ldots\}$, with $T_{n}$ time steps, which may vary across trajectories.
@@ 3.3 Switching Inverse Reinforcement learning
Our SWIRL model is built on the HM-MDP. Instead of explicitly knowing the reward for each mode, we will use IRL to infer these rewards. Mathematically, at each time step $t$, we represent the agent’s internal reward function $r_{z_{t}}$ with an additional dependency on the hidden mode $z_{t}$. This means that the agent receives a reward $r_{z_{t}}$ based on its current hidden mode $z_{t}$, which indicates the decision-making state the animal is in (e.g., water seeking or home seeking), with $r_{z_{t}}$ representing the corresponding intrinsic motivation. Consequently, the optimal policy $\pi_{t}$ is determined by $r_{z_{t}}$. However, SWIRL goes beyond merely embedding IRL within HM-MDP. We also introduce two levels of history dependency into the model. The full graphical model is depicted in Fig. 1.
@@
For action-level history dependency, we treat both the reward and policy under a hidden mode $z$ as functions dependent on the previous $L$ states, specifically $r_{z}:\mathcal{S}^{L}\times\mathcal{A}\rightarrow\mathbb{R}$ and $\pi_{z}:\mathcal{S}^{L}\rightarrow\mathcal{A}$, where $L\in\mathbb{N}$ and $\mathcal{S}^{L}$ denotes the cartesian product of $L$ state spaces. To simplify the notation, we denote an element of $\mathcal{S}^{L}$ as $s^{L}$, so that $r_{z}(s^{L},a):=r_{z}(s^{1},s^{2},\dots,s^{L},a)$ and $\pi_{z}(a|s^{L}):=\pi_{z}(a|s^{1},s^{2},\dots,s^{L})$, and pad both functions with dummy variables if the current time step is less than $L$. We can also add the dependency of previous actions for the reward function. But we use state-only rewards for simplicity and the IRL tradition. It’s straightforward to do so, though. This makes it natural to extend from a single state dependency to a history of state dependencies in our work.
@@ 3.4 SWIRL Inference Procedure
The goal of inference is to learn the hidden modes $z$ and the model parameters $\theta=(\mathcal{P}_{z},r_{z},\pi_{z},p(s_{1}),p(z_{1}))$ given the collected trajectories $\mathcal{D}$. Here, $p(s_{1})$ and $p(z_{1})$ represent the probabilities of the initial state and hidden mode, respectively. The variables $r_{z}$ and $\pi_{z}$ denote the reward and policy associated with the hidden mode $z$, while $\mathcal{P}_{z}$ is the transition matrix between hidden modes. We can maximize the likelihood of the demonstration trajectories $\mathcal{D}$ to learn the optimal $\theta^{*}$, such that $\theta^{*}=\operatorname*{arg\,max}_{\theta}\log{P(\mathcal{D}|\theta)}$ (MLE). However, achieving this objective requires marginalizing over the hidden modes $z$, which is intractable. To address the intractability, we employ the Expectation-Maximization (EM) algorithm, alternating between updating the parameter estimates and inferring the posterior distributions of the hidden modes.

### v2

@@ 1 Introduction
Inverse reinforcement learning (IRL), which infers agents’ policies and intrinsic reward functions based on their interactions with the environment (Ng & Russell, 2000; Abbeel & Ng, 2004; Ziebart et al., 2008, 2010; Wu et al., 2024), has been shown to be effective in capturing animal decision-making intentions by learning reward functions from behavioral trajectories (Sezener et al., 2014; Yamaguchi et al., 2018; Pinsler et al., 2018; Hirakawa et al., 2018). However, traditional IRL assumes a single static reward function over time, limiting its ability to account for shifts in intrinsic motivations. To address this limitation, recent IRL variants have aimed to uncover heterogeneous and time-varying reward functions (Babes-Vroman et al., 2011; Surana & Srivastava, 2014; Nguyen et al., 2015; Ashwood et al., 2022a; Zhu et al., 2024). Despite these advancements, a significant challenge remains unaddressed: animals make decisions based on their history, not just their current state (Kennedy, 2022; Hattori et al., 2019). For example, in perceptual decision-making tasks, mice are found to make new decision based on reward, state and decision history (Ashwood et al., 2022b). Incorporating historical context into the modeling could offer a more accurate representation of animal behavior.
@@
To address the absence of history dependency in time-varying IRL models, we introduce a novel framework called SWitching IRL (SWIRL). Similar to Zhu et al. (2024), SWIRL models long recordings of animal behaviors as a sequence of short-term decision-making processes. Each decision-making process is treated as a Markov decision process (MDP) with a unique reward function that can be inferred using IRL. The segmentation of a long recording into switching decision-making processes is unknown; therefore, each process is regarded as being associated with a hidden mode that must also be inferred. Most importantly, SWIRL incorporates biologically plausible history dependency, drawing on insights from animal behavior. The history dependency is added at two levels: the transitions between decision-making processes (decision-level) and the actions taken to achieve a single goal within a decision-making process (action-level). Decision-level dependency is reflected in the transitions between decision-making processes over extended sequences of time bins, suggesting that an animal’s current choice is shaped by its previous decisions and environmental feedback. Additionally, we posit that these transitions are influenced by the animal’s location. For example, after a mouse drinks from a water port, if it stays nearby, it is more likely to seek another goal. Conversely, if it is far from the port, indicating it has been away for some time, it may become thirsty again and return to search for water. For action-level history dependency, we will model the policy and reward functions as dependent on trajectory history within each decision-making process, using a non-Markovian decision framework. Such a dependency has been studied by existing reinforcement learning research, which often characterizes exploration with reward functions based on historical states and actions (Houthooft et al., 2016; Sharafeldin et al., 2024). Importantly, SWIRL is the first to incorporate history-dependent policies and rewards into IRL.
@@ 2 Related Work
IRL for animal behavior understanding. IRL has been widely used to infer animals’ behavioral strategies and decision-making policies when the reward is unknown. For instance, Pinsler et al. (2018) applies IRL to uncover the unknown reward functions of pigeons, explains and reproduces their flock behavior, and introduces a method to learn a leader-follower hierarchy. Similarly, Hirakawa et al. (2018) uses IRL to learn reward functions from animal trajectories, identifies environmental features preferred by shearwaters, and demonstrates differences in male and female migration route preferences based on the estimated rewards. In another study, Yamaguchi et al. (2018) applies IRL to C. elegans thermotactic behavior, revealing distinct behavioral strategies for fed and unfed worms. Additionally, Sezener et al. (2014) maps reward functions for rats freely moving in a square area, showing how these rewards changed before and after training. While these studies demonstrate the utility of IRL in uncovering behavioral strategies of freely moving animals, they share a key limitation: they all assume a single reward function governs all animal behaviors, which does not account for the complexities of long-term decision-making.
@@
Dynamics-based behavior analysis in animal neuroscience. Traditional approaches to analyzing animal behavior in neuroscience often rely on autoregressive dynamics models. For instance, MoSeq and related works (Wiltschko et al., 2015; Weinreb et al., 2024) assume that animal behavior consists of multiple segments modeled by a hidden markov model (HMM), with each segment evolving through an autoregressive process. Stone (2023) introduces a switching linear dynamical system (SLDS), similar to an ARHMM, but with an additional layer of continuous latent states between the behavioral trajectories and the hidden states representing behavioral segments. We argue that if each segment lasts only a few seconds, it represents meaningful action motifs, such as grooming and sniffing. However, if a segment is significantly longer and reflects a decision-making process, traditional dynamics-based models may not be suitable for identifying these long-term segments. However, these dynamics-based models are not entirely independent of SWIRL. We will demonstrate that SWIRL generalizes purely dynamics-based models by relying on a more principled IRL framework to identify multiple decision-making processes. Our goal is to offer insights that bridge these traditional and new models for animal behavioral analysis.
@@ 3.2 Inverse Reinforcement Learning
Inverse Reinforcement Learning (IRL) addresses the scenario where we have gathered multiple trajectories from an expert agent $\pi^{*}$, comprising a set of state-action pairs $\{(s^{*}_{t},a^{*}_{t})\}$. The goal is to estimate the policy and reward that generated these state-action pairs, often referred to as expert demonstrations in IRL literature. We assume that we have collected $N$ expert trajectories, denoted as $\mathcal{D}=\{\xi_{1},\xi_{2},\ldots,\xi_{N}\}$. Each trajectory consists of a sequence of state-action pairs, represented as $\xi_{n}=\{(s_{1}^{*},a_{1}^{*}),(s_{2}^{*},a_{2}^{*}),\ldots\}$, with $T_{n}$ time steps, which may vary across trajectories.
@@ 3.3 Switching Inverse Reinforcement learning
Our SWIRL model is built on the HM-MDP. Instead of explicitly knowing the reward for each mode, we will use IRL to infer these rewards. Mathematically, at each time step $t$, we represent the agent’s internal reward function $r_{z_{t}}$ with an additional dependency on the hidden mode $z_{t}$. This means that the agent receives a reward $r_{z_{t}}$ based on its current hidden mode $z_{t}$, which indicates the decision-making state the animal is in (e.g., water seeking or home seeking), with $r_{z_{t}}$ representing the corresponding intrinsic motivation. Consequently, the optimal policy $\pi_{t}$ is determined by $r_{z_{t}}$. However, SWIRL goes beyond merely embedding IRL within HM-MDP: We also introduce two levels of history dependency into the model. The full graphical model is depicted in Fig. 1.
@@
For action-level history dependency, we treat both the reward and policy under a hidden mode $z$ as functions dependent on the previous $L$ states, specifically $r_{z}:\mathcal{S}^{L}\times\mathcal{A}\rightarrow\mathbb{R}$ and $\pi_{z}:\mathcal{S}^{L}\rightarrow\mathcal{A}$, where $L\in\mathbb{N}$ and $\mathcal{S}^{L}$ denotes the cartesian product of $L$ state spaces. To simplify the notation, we denote an element of $\mathcal{S}^{L}$ as $s^{L}$, so that $r_{z}(s^{L},a):=r_{z}(s^{1},s^{2},\dots,s^{L},a)$ and $\pi_{z}(a|s^{L}):=\pi_{z}(a|s^{1},s^{2},\dots,s^{L})$, and pad both functions with dummy variables if the current time step is less than $L$. While it is also possible to incorporate previous actions into the reward function, we focus on state-only rewards for simplicity, which is a common practice in IRL. Nonetheless, this framework naturally extends from a single-state dependency to a history of states.
@@ 3.4 SWIRL Inference Procedure
The goal of inference is to learn the hidden modes $z$ and the model parameters $\theta=(\mathcal{P}_{z},r_{z},\pi_{z},p(s_{1}),p(z_{1}))$ given the collected trajectories $\mathcal{D}$. Here, $p(s_{1})$ and $p(z_{1})$ represent the probabilities of the initial state and hidden mode, respectively. The variables $r_{z}$ and $\pi_{z}$ denote the reward and policy associated with the hidden mode $z$, while $\mathcal{P}_{z}$ is the transition matrix between hidden modes. We can maximize the likelihood of the demonstration trajectories $\mathcal{D}$ to learn the optimal $\theta^{*}$, such that $\theta^{*}=\operatorname*{argmax}_{\theta}\log{P(\mathcal{D}|\theta)}$ (MLE). However, achieving this objective requires marginalizing over the hidden modes $z$, which is intractable. To address this issue, we employ the Expectation-Maximization (EM) algorithm, alternating between inferring the posterior distributions of the hidden modes and updating the parameters with the following auxiliary function:
$G(\theta,\hat{\theta})=\log p(s_{1})+\sum_{z}p(z_{1}|\xi,\hat{\theta})\log p(z_{1})$ (1)
$+\sum_{t=1}^{T-1}\log\mathcal{P}(s_{t+1}|s_{t},a_{t})$ (2)
$+\sum_{t=1}^{T}\sum_{z_{t}}p(z_{t}|\xi,\hat{\theta})\log\pi_{z_{t}}(a_{t}|s_{t}^{L};r_{z})$ (3)
$+\sum_{t=1}^{T-1}\sum_{z_{t},z_{t+1}}p(z_{t},z_{t+1}|\xi,\hat{\theta})\log\mathcal{P}_{z}(z_{t+1}|z_{t},s_{t}).$ (4)
Detailed SWIRL algorithm and EM derivation are provided in Appendix A.1A.2.
It is worth noting that this EM derivation seeks to maximize the log-likelihood of policy $\pi_{z_{t}}(a_{t}|s_{t}^{L};r_{z})$ given the posterior of $z_{t}$ (Eq. 3). This procedure can be incorporated into the Maximum Entropy IRL (MaxEnt IRL) (Ziebart et al., 2008, 2010) framework, since optimizing the reward by maximizing $\log\pi$ corresponds to the Lagrangian dual of the standard MaxEnt IRL approach (Zeng et al., 2022). As a result, SWIRL can leverage a variety of IRL methods within the MaxEnt IRL framework: SWIRL can be integrated with model-based methods to infer the policy used to optimize the reward, such as Soft-Q iteration (Haarnoja et al., 2017), or model-free methods, such as Soft-Q learning (Haarnoja et al., 2017) and Soft Actor-Critic (Haarnoja et al., 2018). Alternatively, we can also convert the MaxEnt IRL problem to a model-free inverse Q-learning problem, such as IQ-Learn (Garg et al., 2021). Although model-free methods are often more general and scalable, we find that the model-based Soft-Q iteration approach is more suitable for our behavioral neuroscience study setting, where environment sizes tend to be small, data availability is limited, and the accuracy requirement for reward recovery is high. Accordingly, we use the Soft-Q iteration approach in our main experiments. Nevertheless, in Appendix D.1 we also demonstrate that SWIRL can be integrated with the model-free IQ-Learn to achieve solid performance, illustrating SWIRL’s potential in broader applications.
3.5 Connection to Dynamics-based Behavior Methods
We can show that SWIRL is a more generalized version of ARHMM. In a deterministic MDP, where $p(s_{t+1}|s_{t},a_{t})$ is a delta function and each action $a_{t}$ uniquely determines $s_{t+1}$, $s_{t+1}$ directly implies $a_{t}$. Thus, $\sum_{a_{t}}p(s_{t+1}|s_{t},a_{t})\pi(a_{t}|s_{t},z_{t})=\pi(a_{t}|s_{t},z_{t})=p(s_{t+1}|s_{t},z_{t})$. This effectively reduces SWIRL to ARHMM. In the second real-world experiment (Sec. 4.3), the MDP setup satisfies these assumptions. In such a case, ARHMM can be seen as performing policy learning through behavioral cloning without learning a reward function, whereas SWIRL employs IRL to learn the policy.
Another advanced version of the ARHMM is the switching linear dynamical system (SLDS), which assumes that the state $s_{t}$ is unobserved. Instead, the observed variable $y_{t}$ is a linear transformation of $s_{t}$. Thus, the complete generative model for SLDS consists of: (1) $z_{t}\sim p(z_{t}|z_{t-1})$, (2) $s_{t+1}\sim p(s_{t+1}|s_{t},z_{t})$, and (3) $y_{t+1}\sim p(y_{t+1}|s_{t+1})$. This suggests that the representation $s_{t}$ capturing the primary dynamics is, in fact, a latent representation of the observation $y_{t}$. Building on this concept, we can extend SWIRL into a latent variable model, where $s_{t}$ serves as the latent representation of the true observation state $y_{t}$. This corresponds to the setup of Partial Observation Markov Decision Process (POMDP) in the literature. While SWIRL is not a POMDP method in the formal sense (e.g., modeling belief states or latent dynamics explicitly), it does address non-Markovian behavior by augmenting the state space with recent history (e.g., using ${s_{t},s_{t-1},s_{t-2},…}$). This is a commonly used strategy to handle partial observability or non-Markovianity—similar in spirit to how LSTMs or memory-based policies are used in POMDP settings. Therefore, SWIRL leverages temporal context in a manner similar to methods designed for POMDPs. This extension will link SWIRL to representation learning in reinforcement learning, which can be explored further in future work.
Figure 2: Simulation experiment on a 5 × 5 gridworld. (A) Comparison between the true and discovered reward maps. The color scale represents reward values ranging from 0 to 1. The home reward is defined as $r(s_{t})$, while the water reward depends on both the current and previous locations, $r(s_{t},s_{t-1})$. To present the water reward, each location is divided into five groups, as detailed in the table on the far right. For example, the polygon in the first row represents the reward value when $s_{t}$ is the current location and $s_{t-1}$ is the location to the right. Light grey indicates an impossible transition where no reward exists. (B) Box plots illustrating the Pearson correlation between the true and recovered reward maps, test log-likelihood, and test segmentation accuracy over 10 runs. The x-axis represents the five different models. Outlier selection method is described in Appendix B.8.
4 Results
Throughout the experiment section, we use the following terminology to denote our proposed algorithms and the baseline models we compare:
$\bullet$ MaxEnt (Ziebart et al., 2008, 2010): Maximum Entropy IRL where the reward function is only a function of the current state and action. It is a single-mode IRL approach with a single reward function.
SWIRL code can be found at https://github.com/BRAINML-GT/SWIRL.
4.1 Application to a simulated gridworld environment
To assess the robustness of SWIRL, in Appendix B.7, we evaluated the performance of SWIRL (S-2) under increasing levels of random perturbations in this gridworld dataset. We found that SWIRL can tolerate moderate levels of noise or incomplete data (with 30% of the states and actions in training data permuted), making it suitable for real-world animal behavior datasets where such challenges are common.
4.2 Application to long, non-stereotyped mouse trajectories
Figure 3: Water-restricted labyrinth experiment. (A) Setup for the labyrinth experiment. (B) Boxplot showing held-out test LL over 10 runs, with the x-axis representing the five different models. Outlier selection method is described in Appendix B.8. (C) History dependency inferred by SWIRL (S-2), as reflected in the reward map for the water mode. (D) Inferred reward maps from SWIRL (S-2) under three hidden modes: water, home, and explore. To enhance visualization, the inferred reward $r(s_{t},s_{t-1})$ was averaged over $s_{t-1}$ to produce $r(s_{t})$. (E) Segments of a trajectory from held-out test data, predicted by four SWIRL models. The orange dot indicates when the mouse visits the water port, while the red cross denotes the mouse’s visit to state 0 (home) at that time. (F) Trajectories segmented into hidden modes based on SWIRL (S-2) predictions.
4.2.1 SWIRL inferred interpretable history-dependent reward maps
We applied SWIRL to 80% of the 238 mouse trajectories from the water-restricted labyrinth experiments. According to Rosenberg et al. (2021), mice quickly learned the labyrinth environment and began executing optimal paths from the entrance to the water port within the first hour of the experiment. Therefore, we assume the mice acted optimally concerning the internal reward function guiding their behavior. We tested a number of hidden modes $Z$ from 2 to 5 in Appendix B.6.1 and found $Z=3$ to be the best fit. Fig. 3B displays the held-out test LL for MaxEnt and the SWIRL variations based on the remaining 20% of trajectories. The state dependency in hidden-mode switching dynamics and the history dependency in the reward function contributed to improved test performance. The final SWIRL model (S-2) successfully recovered a water reward map, a home reward map, and an explore reward map (Fig. 3D). For better visualization, we averaged the S-2-recovered history-dependent rewards across previous states and normalized the reward values to a range of (0, 1). In the water reward map, mice received a high reward for visiting the water port. In the home reward map, there was a high reward for visiting state 0 at the center of the labyrinth, which also served as the entrance and exit. Mice occasionally went to state 0 to enter or leave the labyrinth and sometimes passed by on their way to other nodes. In the explore reward map, mice received a high reward for exploring areas of the labyrinth other than state 0 and the water port.
Importantly, SWIRL recovered an interpretable history-dependent reward map for the water port (Fig. 3C). It indicates that mice receive a high reward (1.0) for reaching the water port when their previous location was not the water port. If their prior location was the water port, there is still a reward (0.7) for staying there, but the reward for leaving the water port is even higher (0.9). This observation aligns with the water port design, as mice can only obtain water once every 90 seconds. Consequently, it makes sense that the mice would want to leave the water port after reaching it. Such insights would not be captured by a Markovian reward function that depends solely on the current state.
We then visualized all mouse trajectories based on the hidden-mode segments predicted by SWIRL (S-2) (Fig. 3F). In segments classified as water mode, mice start from various locations in the labyrinth and move toward the water port. In segments identified as home mode, mice begin from distant nodes and head toward the center of the labyrinth (home). In segments categorized as explore mode, mice start from junction nodes or the water port and explore end nodes other than the water port. This result demonstrates that SWIRL can identify sub-trajectories of varying lengths from raw data spanning 500 time points, allowing us to visualize them together and reveal clustered behavioral strategies. This capability has not been achieved by previous studies on freely moving animal behavior over extended recording periods, and we conducted this analysis without prior knowledge of the locations of the water port or home.
We also provide a detailed visualization of the hidden-mode segments from an example trajectory in the held-out test data and compare the segmentation performance of the four SWIRL variations (Fig. 3E). In the S-2 segments, visits to the water port (indicated by orange dots) consistently occur at the end of a water mode segment, while visits to state 0 (home) (indicated by red crosses) typically happen at the conclusion of a home mode segment. Notably, home mode segments that do not include a visit to state 0 can still be valid, as these segments may end at state 1 or 2 (see Appendix C.1). In contrast, the I-1, I-2, and S-1 segments exhibit instances of water segments that do not involve a visit to the water port, along with many home segments that lack clear interpretability. Overall, S-2 successfully identifies robust segments of reasonable length, avoiding the numerous rapid switches seen in the other variations. We attribute this to both the state dependency of hidden mode transitions and the history dependency in rewards. This suggests that mice are unlikely to make quick changes in their decisions; instead, they make choices based on their current location and take into account at least two locations while navigating the maze.
Furthermore, in Appendix B.6.2, we found S-3 and S-4 exhibited even higher test LL than S-2, suggesting the presence of longer action-level history dependency ($L>2$) in this labyrinth dataset. This observation aligns with the partially observable nature of this 127-node labyrinth: The mouse may not know the whole environment, so it likely relies on a longer history of prior states to guide its decision-making. However, hidden mode predictions of S-3 and S-4 are very close to that of S-2, suggesting that S-2 is already powerful enough for hidden mode prediction.
4.3 Application to mouse spontaneous behavior trajectories
Figure 4: Mouse spontaneous behavior experiment. (A) SWIRL (S-1) inferred reward maps for five hidden modes. h1 denotes hidden mode 1, and so on. For better visualization, reward values are normalized to a range of (0, 1). DA corr represents the Pearson correlation between the inferred reward map and the averaged dopamine fluctuation levels. (B) Held-out test LL for each model across different number of hidden modes over 10 runs. The shaded area represents the total area that falls between one standard deviation above and below the mean. (C) Inferred hidden-mode segments for all trajectories, with each row representing a trajectory.
We also employed SWIRL on a dataset in which mice wandered an empty arena without explicit rewards (Markowitz et al., 2023). In this experiment, mouse behaviors were recorded via depth camera video, and dopamine fluctuations in the dorsolateral striatum were monitored. The dataset includes behavior “syllables” inferred by MoSeq (Wiltschko et al., 2015), which indicate the type of behavior exhibited by the mice during specific time periods (e.g., grooming, sniffing, etc.). Consequently, trajectories consist of behavioral syllables, with each time point representing a syllable. We selected 159 trajectories, each comprising 300 time points, by retaining only the 9 most frequent syllables and merging consecutive identical syllables into a single time point. This method, also used in previous reinforcement learning studies on this dataset (Markowitz et al., 2023), ensures that each syllable has sufficient data for learning and allows the model to concentrate on the transitions between different syllables.
The MDP for this experiment comprises 9 states and 9 actions, where the state represents the current syllable and the action signifies the next syllable. As mentioned in Sec. 3.5, the ARHMM can be viewed as a variant of SWIRL that learns the policy through behavior cloning. In other words, the policy for this MDP aligns with the emission probability of the ARHMM. This setup offers an excellent opportunity to compare the performance of SWIRL with ARHMM and its variant, rARHMM.
The best SWIRL model (S-1) recovered reward maps and hidden-mode segments provide insights into the variability of dopamine impacts on animal spontaneous behavior: As illustrated in Fig. 4A, the reward maps exhibit some similarities along with meaningful differences. For certain reward maps, there is a relatively high correlation (e.g., 0.40 and 0.48) with dopamine fluctuations during the corresponding modes. This suggests that dopamine fluctuations can reflect a certain extent of reward during hidden modes 1 and 4.
Furthermore, the plot of hidden mode segments across all trajectories (Fig. 4C) shows how each spontaneous behavior trajectory is temporally segmented into distinct latent modes, each governed by one of the five learned reward functions (h1–h5). Early in the trajectories, we observe a higher prevalence of h1, h2, and h4, while h3 and h5 increasingly dominate toward the later stages. This temporal pattern aligns well with the findings from prior work (Markowitz et al., 2023), which reported that mice exhibit greater activity at the beginning of spontaneous behavior sequences and progressively slow down and disengage over time. The reward maps provide mechanistic insight into these transitions: h1 encourages selective transitions into pause while still supporting other state transitions, representing a modulated rest phase; h2 discourages passivity, with low reward for transitions into pause and strong suppression of grooming; and h4 favors transitions into sniffing, reflecting exploratory drive. In contrast, h3 and h5 strongly reward transitions into pause and suppress departures from it, indicating globally suppressive states that enforce behavioral quiescence. The increasing expression of h3 and h5 over time reflects the mice’s natural progression from active to inactive states, supporting the interpretation that the learned rewards capture underlying behavioral structure consistent with known temporal dynamics.
While the non-Markovian action-level history dependency introduced by SWIRL does not demonstrate superior performance in this particular experiment, the findings showcase SWIRL’s unique contribution to neuroscience research. Specifically, SWIRL serves as a powerful tool for hypothesis testing in behavioral datasets, enabling researchers to validate or challenge hypotheses regarding decision-level dependency as well as non-Markovian action-level dependency. This versatility further confirms SWIRL’s great potential in advancing our understanding of complex behaviors.
5 Discussion
We introduce SWIRL, a novel IRL framework designed to capture history-dependent switching reward functions in complex animal behaviors. Our framework can infer interpretable switching reward functions from lengthy, non-stereotyped behavioral tasks, achieving reasonable hidden-mode segmentation—a feat that, to the best of our knowledge, has not been accomplished previously. Those abilities, combined with its connection to traditional dynamics-based behavior analysis methods in neuroscience, make SWIRL well-suited for behavioral neuroscience applications. Future extensions of SWIRL to partially observable Markov decision processes (POMDPs) could further broaden its utility, offering enhanced insight into complex and nuanced animal behaviors.
Beyond its relevance to neuroscience study, SWIRL also benefits the broader machine learning community. It offers a flexible framework with history-dependent switching reward functions that can be integrated with various model-free IRL algorithms. For example, in Appendix D.1, we highlight SWIRL’s potential to be integrated with IQ-Learn, a state-of-the-art scalable model-free IRL method that has been employed in training large language models (Wulfmeier et al., 2024). Future development of SWIRL that emphasizes scalability, such as incorporating RNNs or transformer-based policy model will be particularly valuable for areas such as robotics and large language models, where adaptable IRL strategies are increasingly in demand.
Impact Statement
SWIRL has the potential to advance neuroscience research by helping uncover connections between complex animal decision-making and corresponding brain activities. In a broader context, SWIRL’s capacity to learn multiple history-dependent switching reward functions can also enable enhanced policy learning and explainability in fields such as robotics and large language models. However, as with all methods that analyze complex behaviors, careful consideration of data privacy, fairness, and responsible deployment is imperative to ensure ethical use.
Acknowledgements
Appendix A Appendix A
A.1 The SWIRL algorithm